OpenStudy (anonymous):

Will medal and fan if helped A fair coin will be tossed three times. What is the probability that two heads and one tail in any order will result? A. 1/2 B. 1/6 C. 2/6 D. 3/8

OpenStudy (anonymous):

C :D

OpenStudy (anonymous):

How tho :/

OpenStudy (anonymous):

The probability of Heads -1/6 The probability of Tails - 1/6 Thus, 2 Head and 1 Tail would be 1/6+ 1/6+1/6 = 3/6 = 1/2 :D

OpenStudy (anonymous):

Oooo

OpenStudy (anonymous):

ooops. The Answer is A. Lol

OpenStudy (welshfella):

nope probability of head in one toss is 1/2

OpenStudy (anonymous):

Isnt is 1/6? since there are 6 attempts?

OpenStudy (welshfella):

i'm just talking about one toss - head or tail

OpenStudy (anonymous):

Oh. okie. :D

OpenStudy (surana):

So it's 1/12?

OpenStudy (welshfella):

possible results are THH or HTH or HHT

OpenStudy (anonymous):

I am confused now... lol

OpenStudy (welshfella):

probability of one of those 3 is 1/2 * 1/2 * 1/2 = = 1/8

OpenStudy (welshfella):

but there are 3 of those - so multiply By 3

OpenStudy (anonymous):

so its 3/8?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

@Dodgegirl is likes.. just give me the answer already ;-; :)

OpenStudy (anonymous):

Oh... Well i totally suck. lol

OpenStudy (anonymous):

nah im just watching being confused XD

OpenStudy (anonymous):

ikr XD

OpenStudy (welshfella):

for the three tosses like HHt the probabilities are multiplied because each toss is independent of the one before

OpenStudy (anonymous):

I need to go back to middle school. lol

OpenStudy (welshfella):

but as HHT THH and HTH are mutually exclusive 9 you cant have both at once) the probabilities are added 1/8 + 1/8 + 1/8 = 3/8